How many atoms are in 56.7 gram mg3p2
1
Expert's answer
2020-12-16T05:53:18-0500

Solution:

Moles of Mg3P2 = Mass of Mg3P2 / Molar mass of Mg3P2

The molar mass of Mg3P2 is 134.88 g mol-1.

Hence,

Moles of Mg3P2 = 56.7 g / 134.88 g mol-1 = 0.4204 mol = 0.42 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of atoms Mg3P2 = 0.42 mol × (6.022×1023 atoms / 1 mol) = 2.53×1023 atoms Mg3P2

Number of atoms Mg3P2= 2.53×1023 atoms


Answer: 2.53×1023 atoms are in 56.7 grams Mg3P2.